I can't figure out how to solve this problem, (5-2)/2x-5 = 3/2x-5
Answer by Nate(3500)

(5-2)/2x-5 = 3/2x-5 orginal
5-2=3 multiple whole equation by 2x-5
3=3 x can be all numbers